Determine the amount of sales (units) that would be necessary under Break-Even Sales Under Present and Proposed Conditi...

  1. Determine the amount of sales (units) that would be necessary under

    Break-Even Sales Under Present and Proposed Conditions

    Darby Company, operating at full capacity, sold 118,800 units at a price of $132 per unit during the current year. Its income statement for the current year is as follows:

    Sales $15,681,600
    Cost of goods sold 7,744,000
    Gross profit $7,937,600
    Expenses:
    Selling expenses $3,872,000
    Administrative expenses 3,872,000
    Total expenses 7,744,000
    Income from operations $193,600

    The division of costs between fixed and variable is as follows:

    Variable Fixed
    Cost of goods sold 70% 30%
    Selling expenses 75% 25%
    Administrative expenses 50% 50%

    Management is considering a plant expansion program that will permit an increase of $1,452,000 in yearly sales. The expansion will increase fixed costs by $145,200, but will not affect the relationship between sales and variable costs.

    Required:

    1. Determine the total variable costs and the total fixed costs for the current year. Enter the final answers rounded to the nearest dollar.

    Total variable costs $
    Total fixed costs $

    2. Determine (a) the unit variable cost and (b) the unit contribution margin for the current year. Enter the final answers rounded to two decimal places.

    Unit variable cost $
    Unit contribution margin $

    3. Compute the break-even sales (units) for the current year. Enter the final answers rounded to the nearest whole number.
    units

    4. Compute the break-even sales (units) under the proposed program for the following year. Enter the final answers rounded to the nearest whole number.
    units

    5. Determine the amount of sales (units) that would be necessary under the proposed program to realize the $193,600 of income from operations that was earned in the current year. Enter the final answers rounded to the nearest whole number.
    units

    6. Determine the maximum income from operations possible with the expanded plant. Enter the final answer rounded to the nearest dollar.
    $

    7. If the proposal is accepted and sales remain at the current level, what will the income or loss from operations be for the following year? Enter the final answer rounded to the nearest dollar.
    $

    8. Based on the data given, would you recommend accepting the proposal?

    1. In favor of the proposal because of the reduction in break-even point.
    2. In favor of the proposal because of the possibility of increasing income from operations.
    3. In favor of the proposal because of the increase in break-even point.
    4. Reject the proposal because if future sales remain at the current level, the income from operations will increase.
    5. Reject the proposal because the sales necessary to maintain the current income from operations would be below the current year sales.

    Choose the correct answer.
Answer #1

1. Total variable costs = 7,744,000*70% + 3,872,000*75% + 3,872,000*50%

= $10,260,800

Total Fixed Costs = 7,744,000*30% + 3,872,000*25% + 3,872,000*50%

= $5,227,200

2.Unit variable cost = total variable cost/number of units ‘

= 10,260,800/118,800

= $86.37 per unit

Unit contribution margin = unit selling price – unit variable cost

= $132-$86.37

= $45.63 per unit

3.Break even units = Total Fixed costs/unit contribution margin

= 5,227,200/45.63

= 114,556.21 units

4.Proposed program = (5,227,200+145,200)/45.63

= 117,738.33 units

5.Desired Income = $193,600

+Fixed Costs 5,372,400

Desired Contribution Margin = $5,566,000

Unit contribution Margin = $45.63

Number of units = 5,566,000/45.63

= 121,981.15 units

6.Maximum Income:

Maximum Sales = 15,681,600+1,452,000= $22,151,700

Contribution Margin = 22,151,700*45.63/132 = $7,657,439.93

Less: Fixed Costs = $5,372,400

Maximum Income = $2,285,039.93

7.Contribution Margin = 15,681,600*45.63/132 = $5,420,844

Less: Fixed Costs = $5,372,400

Income = $48,444

8. b. In favor of the proposal because of the possibility of increasing income from operations.